# The Riemann Hypothesis is false

### v41

@@
$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. For $m\neq n$, let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
@@
We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$. For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=\frac{m}{r}(1+O(r/m))=\frac{m}{r}+O(1)$ hence
@@
Note that if $f(t)$ is a real differentiable function on $[a,b]$, $f^{\prime}(t)$ is monotonic decreasing on $[a,b]$, $f^{\prime}(a)=\beta$ and $f^{\prime}(b)=\alpha,$ then
$\sum_{a<r\leq b}e^{2\pi if(r)}=\sum_{|v|\leq N}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x+o(1)$ (18)
as $N\rightarrow\infty$, where
$\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x\ll\Big{|}\frac{1}{\alpha-v}-\frac{1}{\beta-v}\Big{|}=\frac{\beta-\alpha}{|(\alpha-v)(\beta-v)|}$ (19)
see e.g. [6, pp.10-11]. Taking $f(r)=\frac{t\log(1+\frac{r}{m})}{2\pi},a=0,b=m\varepsilon-1$ and $f^{\prime}(r)=\frac{t}{2\pi(m+r)}$ gives
$\sum_{-N\leq v\leq 0}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $\ll\sum_{-N\leq v\leq 0}\frac{\beta-\alpha}{(\alpha-v)(\beta-v)}$ (20)
$=(\beta-\alpha)\sum_{0\leq v\leq N}\frac{1}{(\alpha+v)(\beta+v)}$ (21)
$\ll(\beta-\alpha)\sum_{0\leq v\leq N}\frac{1}{1+v^{2}}$ (22)
$\ll\beta-\alpha$ (23)
$\sim\frac{t\varepsilon}{2\pi m}.$ (24)
@@
Choose a rational $t\in[T,H]$ so that $\alpha=\frac{t}{2\pi(m+M)}$ and $\beta=\frac{t}{2\pi m}$ are never integers, where $m\in\Big{[}\frac{T}{(1+\varepsilon)\log^{2}T},T(1+\varepsilon)\Big{]}$ is an integer and $M$ is the largest integer $\leq m\varepsilon-1$. Thus $v\neq\alpha$ in particular. Let $K_{0}$ be the largest integer $<\alpha$. We also have
$\sum_{1\leq v\leq\alpha}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ (25)
$\ll\sum_{1\leq v\leq K_{0}}\frac{\beta-\alpha}{(\alpha-v)(\beta-v)}$ (26)
$\ll(\beta-\alpha)\int_{1}^{K_{0}}\frac{1}{(\alpha-v)(\beta-v)}\mathrm{d}v$ (27)
$\ll(\beta-\alpha)\log t$ (28)
$\sim\frac{t\varepsilon}{2\pi m}\log t.$ (29)
Let $K_{1}$ be the least integer $>\beta$. Note that
$\sum_{\beta<v\leq N}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $\ll\sum_{K_{1}\leq v\leq N}\frac{\beta-\alpha}{(v-\alpha)(v-\beta)}$ (30)
$\ll(\beta-\alpha)\int_{K_{1}}^{N}\frac{1}{(v-\alpha)(v-\beta)}\mathrm{d}v$ (31)
$\ll(\beta-\alpha)\log t$ (32)
$\sim\frac{t\varepsilon}{2\pi m}\log t.$ (33)
Since $\beta-\alpha\sim\frac{t\varepsilon}{2\pi m}\ll\frac{\log^{2}T}{T^{1/4}}=o(1)$, note that the range $\alpha<v\leq\beta$ contains no integers. Thus
$\sum_{\alpha<v\leq\beta}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x=0.$ (34)
Combining (34), (33), (29), (24) and (18) yields $\sum_{a<r\leq b}e^{2\pi if(r)}=o(1)$. Infact, by adapting the argument between (18) and (34), one finds that $\sum_{0<r\leq J}e^{2\pi if(r)}=o(1)$ uniformly for $11\leq J\leq b=m\varepsilon-$ 1. Thus by Abel summation it follows that
$\sum_{a<r\leq b}\frac{1}{r}e^{2\pi if(r)}=\sum_{0<r\leq b}\frac{1}{r}e^{it\log(1+\frac{r}{m})}=o(1).$ (35)
Inserting (35) into the right-hand side of (17), summing over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$ and applying the Prime Number Theorem that $\sum_{m\leq x}\Lambda(m)\sim x,$ yields
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (36)
@@
For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$. Thus by arguing as above, one arrives at
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (37)
Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation, invoking (37), (36), (14), (13) and (12) yields
$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (38)
$=-\frac{T}{\pi}\log\log T+o(T)$ (39)
@@
since $H=2T+O(1)$ and $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.182]. By Lemma 2, we have

### v42

@@
$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}$ (9)
say. From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. For $m\neq n$, let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
@@
We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$, where $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=\frac{m}{r}(1+O(r/m))=\frac{m}{r}+O(1)$ hence
@@
Let $J(u)=\sum_{0<r\leq u}e^{2\pi if(r)}=\sum_{0<r\leq M}e^{2\pi if(r)}$ where $u\leq m\varepsilon,M=M(u)$ is the largest integer $\leq u$, $f(r)=\frac{T\log(1+\frac{r}{m})}{2\pi}$ and $m\in\mathbb{N}$. Then $f^{\prime}(r)=\frac{T}{2\pi(r+m)}$. Put $\alpha=f^{\prime}(M),\beta=f^{\prime}(0)$ and choose $T\in\mathbb{Q}$ so that both $\alpha$ and $\beta$ are not integers. Note that $\beta-\alpha\ll\frac{\log^{2}T}{T^{1/4}}=o(1)$. Thus if $\eta$ is any constant $\in(0,\frac{1}{2})$, then the range $[\alpha-\eta,\beta+\eta]$ doesn’t contain any integer. Thus $\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x=0$ and by Lemma 4.7 of Titchmarsh [7] we have
$J(u)=\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x+O(1)\ll 1$ (18)
uniformly for $1\leq u\leq m\varepsilon.$ Then
$\sum_{T^{1/4}<r\leq m\varepsilon-1}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}$ $=\sum_{T^{1/4}<r\leq m\varepsilon-1}\frac{1}{r}e^{2\pi if(r)}$ (19)
[…]
Since $J(u)\ll 1$ uniformly for $1\leq u\leq m\varepsilon,$ it follows from (21) that
$\sum_{T^{1/4}<r\leq m\varepsilon-1}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}\ll T^{-1/4}=\varepsilon$. Applying this into (17) yields
$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}+O(\varepsilon\log m).$ (22)
Putting $\log(1+\frac{r}{m})=\frac{r}{m}+O((r/m)^{2})$ and $e^{O(\frac{r^{2}\log^{4}T}{m})}=1+O(\frac{r^{2}\log^{4}T}{m})$ into the right-hand side of (22) gives
$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log m).$ (23)
Inserting (23) into (17) and summing over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$, yields
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}$ (24)
$+O(T^{3/4}\log T).$ (25)
@@
For $\frac{m}{1+\varepsilon}\leq n\leq m-1,$ put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ so that $\log(n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$. Thus by arguing as above, one arrives at
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{-\frac{irT}{m}}$ (26)
$+O(T^{3/4}\log T).$ (27)
Combining (27) and (25) gives
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (28)
$=\frac{1}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}+O(T^{3/4}\log T).$ (29)
Let $\{{y\}}$ denote the fractional part of $y$. Recall that if $u$ is not an integer, then $\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi(\frac{1}{2}-\{{u\}}+O(1/X))$ as $X\rightarrow\infty$. Hence
$\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}=\sum_{1\leq r\leq T^{1/4}}\frac{2i\sin(rT/m)}{r}=2\pi i\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}.$ (30)
Plugging (30) into (29) gives
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (31)
$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}+O(T^{3/4}\log T).$ (32)
@@
Note that $\sum_{m\leq x}\Lambda(m)\{{x/m\}}=(1-B)x+o(x)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.249]. Also note that $\{{\frac{T}{2\pi m}\}}=\frac{T}{2\pi m}$ for $\frac{T}{2\pi}<m\leq T(1+\varepsilon)$. Thus
$\sum_{m\leq\frac{T}{(1+\varepsilon)\log^{2}T}}\Lambda(m)\{{T/2\pi m\}}-\sum_{m\leq\frac{T}{2\pi}}\Lambda(m)\{{T/2\pi m\}}-\sum_{\frac{T}{2\pi}<m\leq T(1+\varepsilon)}\Lambda(m)\{{T/2\pi m\}}$ (33)
$+\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}+O(T^{-1/4})\Big{)}$ (34)
$=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}\sum_{\frac{T}{2\pi}<m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{m}+\frac{T}{2}+o(T)$ (35)
$=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}(\log T-\log(T/2\pi))+\frac{T}{2}+o(T)$ (36)
$=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}\log(2\pi)+\frac{T}{2}+o(T)$ (37)
since $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.187]. Plugging (37) into (32) gives
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (38)
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}\log(2\pi)+\frac{T}{2}+o(T)\Big{)}.$ (39)
By a similar argument, one arrives at
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}$ (40)
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}-\frac{H}{2\pi}\log(2\pi)+\frac{T}{2}+o(T)\Big{)}$ (41)
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{\pi}-\frac{T}{\pi}\log(2\pi)+\frac{T}{2}+o(T)\Big{)}$ (42)
since $H=2T+O(1)$. Combining (42) and (39) gives
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (43)
$-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}$ (44)
$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(2\pi)+o(T)\Big{)}.$ (45)
Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation, invoking (45), (14), (13) and (12) yields
$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(2\pi)+o(T)$ (46)
$=-\frac{T}{\pi}\log\log T+(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(2\pi)+o(T).$ (47)
By Lemma 2, we have